determine between which consecutive intergers that real zeros of f(x)=x^3-2 are located
Answer by JBarnum(2146)

x cant be a negative number or it would be negative + negative which is less than 0
x cant be 2 or more because 2^3= 8 8-2=6 which isnt 0
x cant be 1 or less because 1^3=1 and 1-2 =-1 which isnt 0
so the number has to be between the two consecutive intergers: 1 and 2


and to be exact if your wondering whats exactly x to make the zero it is 1.25992104989487316476...